OpenStudy (anonymous):

What is the y−coordinate of the y−intercept of the line that passes through the points (−4, −4) and (4, 8)?

OpenStudy (anonymous):

Slope is rise/run = change in y/change in x = (y2 - y1)/(x2 - x1) = (8 - -4)/(4 - -4) = 12/8 = 3/2. Since it's a line, we can represent it as y = mx + b where m and b are constants. We just found m, and we have some values of x and y to test out. Plugging everything in: -4 = (3/2)*-4 + b -4 = -6 + b b = 2 y = (3/2)x + 2 Y intercept is when x = 0, and when you plug 0 in for x all that is left is the constant, b, which is 2 in this case. Answer is 2.

OpenStudy (anonymous):

the answer should be 2

OpenStudy (anonymous):

Thanks a bunch, Blue! I really appreciate you explaining too. GO BLUE PIG. WOO. ^-^